OpenStudy (anonymous):

convert the angle 150 into radian s

OpenStudy (precal):

multiply it by pi/180

OpenStudy (precal):

\[150\left( \frac{ \pi }{ 180 } \right)=\frac{ 150 }{ 180 }\pi=\frac{ 5 }{ 6 }\pi\]

OpenStudy (xapproachesinfinity):

Correct